# Second version: August 2024 Third version: November 2024 This version: September 2025)

### v2

@@ 2.2 Test statistics, size controllability, and a new result
It is now easy to see that $\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$ and that $\mathsf{B}$ is a linear space (cf. also Lemma 3.1 in Pötscher and Preinerstorfer (2025)). Simple examples can be constructed to show that $\mathop{\mathrm{s}pan}(X)\neq\mathsf{B}$, in general, cf. Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) as well as Examples A.1-A.4 in Appendix A further below.
@@
Note that condition (10) can equivalently be expressed in terms of certain diagonal elements of the ‘hat matrix’ $H$, see (12) in Remark 2.1 below. This is, however, not the case for condition (7). An informal verbal description of (10) is given in (1) in the Introduction.
@@
Most importantly, the above theorem shows that, given Assumption 1, the condition in (7) is sufficient for the existence of a (finite) size-controlling critical value $C(\alpha)$ satisfying (6), while the weaker condition (10) is necessary. Furthermore, in case the design matrix $X$ and the vector $R$ are such that $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$, and hence the condition in (7) coincides with that in (10), the condition (7) is also necessary. However, $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ is not always true (see Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) or the examples in Appendix A further below), although the equality holds generically (cf. Theorem 3.9 and Lemma A.3 in Pötscher and Preinerstorfer (2018)). We now show in the subsequent theorem that in the situation considered in this article, namely testing only a single restriction, the condition in (7) in Theorem 2.1 can actually be replaced by that in (10).
Theorem 2.2.
  1. 1.
    For every $0<\alpha<1$ there exists a real number $C(\alpha)$ such that (6) holds, provided that (10) holds. Furthermore, under condition (10), even equality can be achieved in (6) by a proper choice of $C(\alpha)$, provided $\alpha\in(0,\alpha^{\ast}]\cap(0,1)$ holds, where $\alpha^{\ast}$ given by (8) is positive and where $C^{\ast}$is given by (9) for $\mu_{0}\in\mathfrak{M}_{0}$ (with neither $\alpha^{\ast}$ nor $C^{\ast}$ depending on the choice of $\mu_{0}\in\mathfrak{M}_{0}$).
  2. 2.
  3. 3.
  4. 4.
    If (10) is violated, then $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every choice of critical value $C$, every $\mu_{0}\in\mathfrak{M}_{0}$, and every $\sigma^{2}\in(0,\infty)$ (implying that size equals $1$ for every CC). [^9] Cf. Footnote 8.
The main take-away of Theorem 2.2 is that given Assumption 1 holds, the condition in (10) is necessary and sufficient for the existence of a (smallest) finite size-controlling critical value when one is testing only a single restriction. Note that the conditions in (7) and (10) do not depend on the weights used in the construction of the covariance matrix estimator or on $r$. They only depend on $X$ and $R$. This and more (e.g., how the conditions relate to high-leverage points) is discussed subsequent to Theorem 5.1 (and in Remarks 5.2-5.4, 5.6, and 5.9) in Pötscher and Preinerstorfer (2025) to which we refer the reader for a detailed account. As a point of interest we also note that condition (10) given above is exactly the same as condition (8) in Pötscher and Preinerstorfer (2025) (with $q=1$); in that reference, the latter condition is shown to be necessary and sufficient for size control of the standard (uncorrected) F-test statistic (regardless of whether $q=1$ or not).
@@
(iii) The condition $e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)$ for every $i=1,\ldots,n$” (which is tantamount to $h_{ii}<1$ for every $i=1,\ldots,n$”) implies (10), and thus is sufficient for size-controllability of $T_{Het}$ (but not necessary, see, e.g., Example A.2).
@@ 3 Conclusion
We finally remark that the algorithms designed to compute size-controlling critical values as discussed in Section 10 and Appendix E of Pötscher and Preinerstorfer (2025) can be used as they stand also in situations where (a single restriction is tested and) size controllability has been verified through checking condition (10) and appealing to Theorem 2.2, but where (7) does not hold. This is so since the discussion of the before mentioned algorithms in Pötscher and Preinerstorfer (2025) only requires existence of a (finite) size-controlling critical value, but does not depend on the way this existence is verified.
Appendix A Auxiliary results

### v3

# Second version: August 2024 Third version: November 2024 Fourth version: September 2025

@@ 2.2 Test statistics, size controllability, and a new result
It is now easy to see that $\mathop{\mathrm{s}pan}(X)\subseteq\mathsf{B}$ and that $\mathsf{B}$ is a linear space (cf. also Lemma 3.1 in Pötscher and Preinerstorfer (2025)). Simple examples can be constructed to show that $\mathop{\mathrm{s}pan}(X)\neq\mathsf{B}$, in general; cf. Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) as well as Examples A.1-A.4 in Appendix A further below.
@@
Most importantly, the above theorem shows that, given Assumption 1, the condition in (7) is sufficient for the existence of a (finite) size-controlling critical value $C(\alpha)$ satisfying (6), while the weaker condition (10) is necessary. Furthermore, in case the design matrix $X$ and the vector $R$ are such that $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$, and hence the condition in (7) coincides with that in (10), the condition (7) is also necessary. However, $\mathsf{B}=\mathop{\mathrm{s}pan}(X)$ is not always true (see Example C.1 in Appendix C of Pötscher and Preinerstorfer (2025) or the examples in Appendix A further below), although the equality holds generically (cf. Theorem 3.9 and Lemma A.3 in Pötscher and Preinerstorfer (2018)). We now show in the subsequent theorem that in the situation considered in this article, namely testing only a single restriction, the condition in (7) in Theorem 2.1 can actually always be replaced by that in (10). Before we present that theorem, we discuss an equivalent formulation of condition (10) that is expressed in terms of certain diagonal elements of the ‘hat matrix’ $H$, see (12) below. [^9] An informal verbal description of (10) is given in (1) in the Introduction.
@@
We now present the announced theorem.
Theorem 2.2.
 [^12] Following the suggestion of some readers, we mention the equivalent condition (12) explicitly in this theorem, although the equivalence with (10) has already been noted in Remark 2.1.
  1. 1.
    For every $0<\alpha<1$ there exists a real number $C(\alpha)$ such that (6) holds, provided that (10) (or equivalently (12)) holds. Furthermore, under condition (10) (or equivalently (12)), even equality can be achieved in (6) by a proper choice of $C(\alpha)$, provided $\alpha\in(0,\alpha^{\ast}]\cap(0,1)$ holds, where $\alpha^{\ast}$ given by (8) is positive and where $C^{\ast}$is given by (9) for $\mu_{0}\in\mathfrak{M}_{0}$ (with neither $\alpha^{\ast}$ nor $C^{\ast}$ depending on the choice of $\mu_{0}\in\mathfrak{M}_{0}$).
  2. 2.
    Suppose (10) (or equivalently (12)) is satisfied. Then a smallest critical value, denoted by $C_{\Diamond}(\alpha)$, satisfying (6) exists for every $0<\alpha<1$. And $C_{\Diamond}(\alpha)$ is also the smallest among the critical values leading to equality in (6) whenever such critical values exist.
  3. 3.
    Suppose (10) (or equivalently (12)) is satisfied. Then any $C(\alpha)$ satisfying (6) necessarily has to satisfy $C(\alpha)\geq C^{\ast}$. In fact, for any $C<C^{\ast}$ we have $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every $\mu_{0}\in\mathfrak{M}_{0}$ and every $\sigma^{2}\in(0,\infty)$.
  4. 4.
    If (10) (or equivalently (12)) is violated, then $\sup_{\Sigma\in\mathfrak{C}_{Het}}P_{\mu_{0},\sigma^{2}\Sigma}(T_{Het}\geq C)=1$ for every choice of critical value $C$, every $\mu_{0}\in\mathfrak{M}_{0}$, and every $\sigma^{2}\in(0,\infty)$ (implying that size equals $1$ for every CC). [^13] Cf. Footnote 8.
The main take-away of Theorem 2.2 is that, given Assumption 1 holds, the condition in (10) (or equivalently (12)) is necessary and sufficient for the existence of a (smallest) finite size-controlling critical value when one is testing only a single restriction. [^14] By contraposition, the design matrices $X$ and restrictions $R$ for which size control fails are precisely characterized by failure of (10) (or equivalently (12)). One example is when $X$ contains the dummy $e_{i}(n)$ as its first column, say, and $R=(1,0,\ldots,0)$ (or, more generally, $R$ has a non-zero first entry). Another example arises when the first two columns of $X$ are given by $(1,\ldots,1)^{\prime}$ and $(1,-1,\ldots,-1)^{\prime}$, and the first two entries of $R$ are both equal to $1$. The condition $e_{i}(n)\notin\mathop{\mathrm{s}pan}(X)$ for every $i=1,\ldots,n$” (which is tantamount to $h_{ii}<1$ for every $i=1,\ldots,n$”) implies (10), and thus is sufficient for size-controllability of $T_{Het}$ (but not necessary, see, e.g., Example A.2). Note that the conditions in (7), (10), as well as (12) do not depend on the weights used in the construction of the covariance matrix estimator or on $r$. They only depend on $X$ and $R$. This and more (e.g., how the conditions relate to high-leverage points) is discussed subsequent to Theorem 5.1 (and in Remarks 5.2-5.4, 5.6, and 5.9) in Pötscher and Preinerstorfer (2025) to which we refer the reader for a detailed account. As a point of interest we also note that condition (10) given above is exactly the same as condition (8) in Pötscher and Preinerstorfer (2025) (with $q=1$); in that reference, the latter condition is shown to be necessary and sufficient for size control of the standard (uncorrected) F-test statistic (regardless of whether $q=1$ or not).
@@ 3 Conclusion
We finally remark that the algorithms designed to compute size-controlling critical values as discussed in Section 10 and Appendix E of Pötscher and Preinerstorfer (2025) can be used as they stand also in situations where (a single restriction is tested and) size controllability has been verified through checking condition (10) (or equivalently (12)) and appealing to Theorem 2.2, but where (7) does not hold. This is so since the discussion of the before mentioned algorithms in Pötscher and Preinerstorfer (2025) only requires existence of a (finite) size-controlling critical value, but does not depend on the way this existence is verified.
Appendix A Auxiliary results and some examples